OpenStudy (anonymous):

Plz help :) easy medal Solve for x 16x2 − 81 = 0.

OpenStudy (anonymous):

A. x = 4, 9 B. x = −4, −9 C. x = 9 over 4, x = −9 over 4 D. x = 81 over 16, x = −81 over 16

Nnesha (nnesha):

difference of square \[\huge\rm a^2- b^2= (a+b)(a-b)\] take square root of both terms write your answer in two parentheses (square root of first term + sqrt of 2nd term) (sqrt of 1st term - sqrt of 2nd term)

Nnesha (nnesha):

and then set both parentheses equal to zero solve for x!

OpenStudy (anonymous):

ty for your help :) @Nnesha

Nnesha (nnesha):

my pleasure